# Feedback Delay Network Optimization

### v1

@@ 3.5 Loss function
Figure 5: Evolution of losses (left) and the relative density increase of the IR (right).
@@
$\mathcal{L}=\mathcal{L}_{\textrm{spectral}}(\bm{H}(\bm{z}_{\mu}))+\mathcal{L}_{\textrm{sparsity}}(\bm{U})$ (18)
and the individual loss terms are
$\mathcal{L}_{\textrm{spectral}}(\bm{H}(\bm{z}_{\mu}))$ $=\frac{1}{\mu}\sum_{z\in\bm{z}_{\mu}}\sum_{i=1}^{N}\left({\lvert H_{i}(z)\rvert-1}\right)^{2}$
@@
$\mathcal{L}_{\textrm{sparsity}}(\bm{U})$ $=\frac{N\sqrt{N}-\sum_{i,j}\lvert U_{ij}\rvert}{N(\sqrt{N}-1)}$
where $H_{i}(z)$ is the output of the network’s $i^{\text{th}}$ channel computed from the output of the $i^{\text{th}}$ delay line and scaled by $c_{i}$.
The sparsity loss, which is the last term above, is calculated from the sum of the absolute values of the matrix entries, i.e. $\sum_{i,j}\lvert U_{ij}\rvert$. For orthogonal matrices, its value is bounded by $-N\sqrt{N}$ and $N$. To make the dynamic range of the $\mathcal{L}_{\textrm{sparsity}}$ comparable with that of $\mathcal{L}_{\textrm{spectral}}$ we shifted $\sum_{i,j}\lvert U_{ij}\rvert$ by $N$ and normalized it by $N(\sqrt{N}-1)$. The term is then shifted by $1$ to span positive values only.
In our previous work [9], the sparseness was controlled by a temporal loss consisting in the ratio of the $\ell_{1}$ norm to the $\ell_{2}$ norm of the estimated IR. This required the inverse Fourier transform, slowing down the optimization steps.
To visualize the impact the loss function (18) has on the IR density, Fig. 5 illustrates how the $\ell_{1}$ norm of the estimated IR signal $h$, i.e., $\sum_{k}\lvert h(k)\rvert$, evolves in relation to the losses at each epoch of one run of training. The right pane shows the increase in density relative to initialization. The left pane shows the evolution of the loss and its terms at each epoch. The terms decay at similar rates, and after 10 epochs, they start to converge. Compared to the losses, the increase in density starts to converge quicker, reaching a $16\%$ increase after just 5 epochs. Without sparsity control, the magnitude response may converge towards that of a comb filter, leading to a corresponding decrease in IR density, as demonstrated in our previous work [9].
@@ 4.1 Frequency-dependent decay control
The attenuation filter is designed to approximate a target frequency-dependent reverberation time $T_{60}(\omega)$ by achieving a target gain-per-sample $\gamma(\omega)$
@@
Attenuation filters are commonly introduced after the delay lines [20]. To simplify the transfer function, at the price of a negligibly lower $T_{60}(\omega)$ accuracy, we place the filters in the feedback loop, resulting in the following transfer function
@@ 4.3 Optimizing Householder matrix
When using the Householder matrix $\bm{U}=\bm{I}-2\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}$, where $\bm{v}$ is a unit vector, gives a simple, yet restricted, parametrization of orthogonal matrices. Using the Sherman-Morrison formula, i.e.,
(${\lparen\bm{W}+2\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}\rparen}^{\text{-}1}={\bm{W}}^{\text{-}1}-\frac{2{\bm{W}}^{\text{-}1}\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}{\bm{W}}^{\text{-}1}}{1+2\bm{v}^{\mathstrut\scriptscriptstyle\top}{\bm{W}}^{\text{-}1}\bm{v}}$ (25)
@@
$H(z)=\bm{c}^{\top}{\left({\bm{W}(z)+2\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}}\right)}^{\text{-}1}\bm{b}\,,$ (26)
where $\bm{W}(z)={\bm{D_{\bm{m}}(z)}}^{\text{-}1}-\bm{I}$. With this parameterization, the computational complexity required to calculate the FDN transfer function at one frequency point is reduced from $\mathcal{O}(N^{3})$ to $\mathcal{O}(N)$.
@@
$H(z)=\bm{c}^{\top}{\left({\bm{W_{\gamma}}(z)+2\bm{v}\bm{v_{\gamma}^{\mathstrut\scriptscriptstyle\top}}}\right)}^{\text{-}1}\bm{b}\,,$ (27)
where $\bm{W_{\gamma}}(z)={\bm{D_{\bm{m}}(z)}}^{\text{-}1}-\bm{\Gamma}$ and $\bm{v_{\gamma}}^{\mathstrut\scriptscriptstyle\top}=\bm{v}^{\mathstrut\scriptscriptstyle\top}\Gamma$. With this parametrization, the optimizable parameters become $\bm{v},\,\bm{b},\,\text{and}\,\bm{c}$. The input to the sparsity loss remains $\bm{U}$, which is now defined using the Householder transformation.
@@ 5.1 Evaluation setup
We concentrate on smaller FDNs with limited delay lines, as these are significantly impacted by coloration. Specifically, we consider FDN sizes of $N=4,6,8$ The values of the delay-line lengths are presented in Table 1. The delay lengths are prime numbers distributed logarithmically. In all configurations, the total number of modes is $90008700<\mathcal{M}<$ 9000.
Table 1: Delay-line lengths for each size $N$ of the analyzed FDNs. Values are logarithmically distributed prime numbers.
@@
During optimization, we used a sampling rate of $f_{\textrm{s}}=48$ kHz and $M=480000$ frequency points equally distributed over $[0,\pi]$. The batch size was $\mu=2000$, requiring 240 steps to complete an epoch. For training, we used $80\%$ of the frequency points, while the remaining $20\%$ were used for validation. We employed Adam optimizer [27] with learning rate of $\eta=10^{-3}$. Regarding the gain-per-sample, we found that $\gamma=0.9999$ leads to stable training and fast convergence. This choice implies $T_{\textrm{60}}=1.439\,$s and satisfies (17).
The configurations under analysis include the differentiable FDN optimized according to our prior work (DAFx23) [9], the one optimized with the proposed method (DiffFDN), and two variations of the latter: one with an optimizable Householder matrix (HH) and another with an optimizable scattering matrix (SCAT).
@@ 5.2 Computational complexity
The presented optimization utilizes a sparse set of frequency sampling points at each step. This choice is driven by the aim to decrease the overall number of operations, providing a notable reduction when compared to the baseline method [9]. During the forward pass, the computational complexity needed to calculate the FDN transfer function at a single frequency point in (13) is $\mathcal{O}(N^{3})$. For each epoch, the proposed method (DiffFDN) computes the transfer function at $M$ points. In contrast, the baseline approach (DAFx23) requires a similar number of frequency points at each training step, resulting in a linear increase in the number of operations with the dataset size, e.g., 256 in [9]. This is because the number of epochs required for convergence in the baseline is close to that of the presented method. Additionally, in the forward pass of the baseline, the FFT is necessary due to the temporal loss, introducing additional operations. By employing the Householder matrix, we can significantly decrease the computational complexity of (13) to $\mathcal{O}(N)$. Introducing the FFT would raise this complexity to $\mathcal{O}(N\text{log}_{2}(N))$.
@@
DiffFDN 208 324 448
DiffFDN (HH) 200 300 400
DiffFDN (SCAT) 288 480 704
@@ 5.3 Modal excitation distribution
Table 3: Standard deviation of the modal excitation before (Init) and after optimization (Optim) at different FDN sizes.
@@
To ensure that the changes to the optimization framework did not reduce the performance of the method introduced in [9], we conducted identical tests using our previous optimization framework. Each set of parameters used at initialization is shared between the two methods. We observed comparable convergence rates for both optimization methods. Consequently, we allowed each training run to complete 20 epochs before halting it and extracting the output parameters for evaluation.
@@
Table 4 also shows the results of a similar test run on the FDN with optimizable scattering matrix (SCAT) and Householder matrix (HH), both of size $N=4$. Due to a different parameterization of the feedback matrix, the distributions of the initial parameters are different from the standard configuration. After optimization, the standard deviation reduces in both cases. Among all tested configurations, the HH case exhibited the smallest improvement, likely due to its more restrictive parameterization. The SCAT configuration also showed a lower degree of improvement compared to the standard configuration. This could be attributed to the higher complexity of the problem, suggesting potential enhancement through fine-tuning the optimization framework.
Table 4: Standard deviation of the modal excitation before (Init) and after optimization (Optim) of different frameworks, each employing an FDN of size N=4 and identical set of dealys.
@@
DiffFDN 7.8346 4.4518
DiffFDN (HH) 8.5641 7.8765
DiffFDN (SCAT) 7.6200 5.8822
@@ 6.1 Listening test procedure
Figure 6: Listening test results on the analysized reverberation conditions: (a) lossless, (b) homogeneous decay, and (c) frequency-dependent decay.
@@
In the part testing the lossless condition, we evaluated late non-decaying reverb. The reference for this part was a 3-second segment of white Gaussian noise, known to be highly colorless. The IRs of the FDNs were extracted from the mixing time, thereby eliminating the initial echo buildup from the evaluation. The section testing homogeneous decay, focused on frequency-independenty decaying IRs, with a $T_{60}$ value set at 2 seconds for all frequencies. A corresponding exponential decaying envelope with the same $T_{60}$ value was applied to the reference white Gaussian noise. The part testing frequency-dependent decay involved decaying IRs with a fixed frequency dependent $T_{60}(\omega)$ across conditions. The reference in this case was the IRs of an FDN of size $N=64$, ensuring a sufficiently smooth response. The frequency-dependent attenuation filter was designed following the steps outlined in Section 4.1.
The test evaluated the coloration of the standard differentiable FDN (DiffFDN) proposed in this paper, the DiffFDN with scattering matrix (SCAT), and that with optimizable Householder matrix (HH). Each configuration was tested on a separate page where the number and lengths of the delays were constant, and only the feedback matrix and the input and output gains were altered. The FDN implementation of the Schroeder series allpass reverberator (SH) was the anchor, whereas the random orthogonal feedback matrix (RO) acted as a baseline. For the RO condition, the initial values of optimization of the DiffFDN were used. The direct gain $d$ was set to zero in all cases. In total, there were 18 listening test pages with six stimuli each. The loudness of each IR was normalized to ensure consistency across conditions.
@@
The $p$-values for most paired conditions indicate significant differences among all pairs of results. However, there are a few exceptions observed in the distribution of DiffFDN and HH. In these cases, the $p$-values exceed the alpha level across all conditions with decaying modes, suggesting a convergence in the scores assigned to the two conditions. Moreover, DiffFDN and HH reject the null hypothesis when paired with SCAT in the lossy cases for $N=8$, suggesting that for larger FDN sizes, the performance improvement provided by SCAT diminishes.
The IRs of the optimized FDNs always show a median greater than the initial condition (RO), proving the effectiveness of the optimization. In the late reverb analysis, Fig. 6 (a), the DiffFDN has median ratings of 74, 82.5, and 70, respectively, for increasing FDN size. Only for $N=$ 8, the FDN with scattering matrix received ratings higher than DiffFDN with a median of 76.
Introducing early reflections provides an advantage for the FDN with a scattering matrix (SCAT) compared to other conditions. In general, spectral flatness is associated with temporal smoothness. The scattering feedback matrix increases echo density during early reflections, potentially suggesting a less colored response, as shown in Fig. 6 (b) and (c). This effect becomes more noticeable for smaller FDNs, and for $N=8$, the medians of the results begin to converge among optimized FDNs.
The confidence intervals for the results of all tests are relatively large. Although the objective results in Table 4 indicate a limited reduction in modal excitation standard deviation, the optimized Householder matrix was rated relatively close to the DiffFDN in the listening test. The distribution of modal excitation is one quantity that can aid in deducing the coloration of the reverberator. However, this is a much more complicated concept, as evidenced by the results presented in the listening test.

### v2

@@ 3.5 Loss function
Figure 5: Evolution of losses (left) during training of a size 8 FDN with an orthogonal feedback matrix trained with $\alpha=1$ (blue) and $\alpha=0$ (green). On the right, the echo density profile [21] at initialization and after optimization for both cases.
@@
$\mathcal{L}=\mathcal{L}_{\textrm{spectral}}(\bm{H}(\bm{z}_{\mu}))+\alpha\mathcal{L}_{\textrm{sparsity}}(\bm{U})$ (16)
where $\alpha$ is a weighting term that controls the influence of $\mathcal{L}_{\textrm{sparsity}}$ on the overall loss. In this work, we set $\alpha$ to 1. The individual loss terms are
$\mathcal{L}_{\textrm{spectral}}(\bm{H}(\bm{z}_{\mu}))$ $=\frac{1}{\mu}\sum_{z\in\bm{z}_{\mu}}\frac{1}{N}\sum_{i=1}^{N}\left({\lvert H_{i}(z)\rvert-1}\right)^{2}$ (17)
@@
$\mathcal{L}_{\textrm{sparsity}}(\bm{U})$ $=\frac{N\sqrt{N}-\sum_{i,j}\lvert U_{ij}\rvert}{N(\sqrt{N}-1)},$ (18)
where $H_{i}(z)$ is the output of the network’s $i^{\text{th}}$ channel computed from the output of the $i^{\text{th}}$ delay line and scaled by $c_{i}$. Including the channel-wise loss term in (17) is optional, as our tests showed comparable results on the objective metrics with and without it.
The sparsity loss, which is the last term above, is calculated from the sum of the absolute values of the matrix entries, i.e. $\sum_{i,j}\lvert U_{ij}\rvert$. For orthogonal matrices, its value is bounded between $N$ and $N\sqrt{N}$. To make the dynamic range of the $\mathcal{L}_{\textrm{sparsity}}$ comparable with that of $\mathcal{L}_{\textrm{spectral}}$ we shifted $\sum_{i,j}\lvert U_{ij}\rvert$ by $N\sqrt{N}$ and normalized it by $N(\sqrt{N}-1)$. This sparsity loss is motivated by the fact that echo density increases when the delay lines are highly interconnected. This principle underlies the FDN’s structure, distinguishing it from that of parallel comb-filters [14, 3]. Moreover, in our previous work [9] we found that the absence of a term controlling the temporal density of the system’s IR may lead the matrix $\bm{U}$ to converge towards either a diagonal matrix or its permutation. In our previous work [9], the sparseness was controlled by a temporal loss consisting in the ratio of the $\ell_{1}$ norm to the $\ell_{2}$ norm of the estimated IR. This required the inverse DFT, slowing down the optimization steps.
To clarify the impact of the loss function (16) on echo density, Fig. 5 illustrates how echo density profile, as defined in [21], is influenced by the proposed training, particularly by the inclusion of $\mathcal{L}_{\textrm{sparsity}}$. The left panel in Fig. 5 displays the evolution of the loss and its components over each epoch. The plot shows the losses when an instance of an FDN with eight delay lines is trained using (16) with $\alpha=1$, as well as using only $\mathcal{L}_{\textrm{spectral}}$, i.e. $\alpha=0$ in (16), indicated in the plot with an asterisk.
In Fig. 5 (left), the terms decay at similar rates, and after six epochs, they begin to converge. The right panel presents the echo density of the FDN at initialization, and after optimization with $\alpha=1$ and $\alpha=0$. Both optimizations demonstrate an improvement in echo density, as the curve reaches 1 faster. However, when $\mathcal{L}_{\textrm{sparsity}}$ is included (i.e., $\alpha=1$), the echo buildup occurs more rapidly from the early reflections. Moreover, without sparsity control, the magnitude response may converge towards that of a comb filter, leading to a corresponding decrease in IR density, as demonstrated in our previous work [9].
To implement the differentiable model and the training iterations, we used the PyTorch library [22]. PyTorch handles parameter adjustments through backpropagation, which involves computing the gradient of the loss function with respect to the given parameters. PyTorch’s built-in differentiation engine automatically computes the gradients for any computational graph. Additionally, PyTorch supports complex-valued tensors and backpropagation for real-valued functions of complex tensors, using Wirtinger calculus, which is particularly beneficial for our work in the frequency domain [23].
@@ 4.1 Frequency-dependent decay control
The attenuation filter is designed to approximate a target frequency-dependent reverberation time $T_{60}(\omega)$ by achieving a target gain-per-sample $\gamma(\omega)$:
@@
Attenuation filters are commonly introduced after the delay lines [24]. To simplify the transfer function, at the price of a negligibly lower $T_{60}(\omega)$ accuracy, we place the filters in the feedback loop, resulting in the following transfer function:
@@ 4.3 Optimizing Householder matrix
When using the Householder matrix $\bm{U}=\bm{I}-2\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}$, where $\bm{v}$ is a unit vector, gives a simple, yet restricted, parametrization of orthogonal matrices. Using the Sherman-Morrison formula for an invertible square matrix $\bm{V}$, i.e.,
(${\lparen\bm{V}+2\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}\rparen}^{\text{-}1}={\bm{V}}^{\text{-}1}-\frac{2{\bm{V}}^{\text{-}1}\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}{\bm{V}}^{\text{-}1}}{1+2\bm{v}^{\mathstrut\scriptscriptstyle\top}{\bm{V}}^{\text{-}1}\bm{v}}$ (25)
@@
$H(z)=\bm{c}^{\top}{\left({\bm{V}(z)+2\bm{v}\bm{v}^{\mathstrut\scriptscriptstyle\top}}\right)}^{\text{-}1}\bm{b}\,,$ (26)
where $\bm{V}(z)={\bm{D_{\bm{m}}(z)}}^{\text{-}1}-\bm{I}$. With this parameterization, the computational complexity required to calculate the FDN transfer function at one frequency point is reduced from $\mathcal{O}(N^{3})$ to $\mathcal{O}(N)$.
@@
$H(z)=\bm{c}^{\top}{\left({\bm{V_{\gamma}}(z)+2\bm{v}\bm{v_{\gamma}^{\mathstrut\scriptscriptstyle\top}}}\right)}^{\text{-}1}\bm{b}\,,$ (27)
where $\bm{V_{\gamma}}(z)={\bm{D_{\bm{m}}(z)}}^{\text{-}1}-\bm{\Gamma}$ and $\bm{v_{\gamma}}^{\mathstrut\scriptscriptstyle\top}=\bm{v}^{\mathstrut\scriptscriptstyle\top}\Gamma$. With this parametrization, the optimizable parameters become $\bm{v},\,\bm{b},\,\text{and}\,\bm{c}$. The input to the sparsity loss remains $\bm{U}$, which is now defined using the Householder transformation.
@@ 5.1 Evaluation setup
We concentrate on smaller FDNs with limited delay lines, as these are significantly impacted by coloration. Specifically, we consider FDN sizes of $N=4,6,8$ The values of the delay-line lengths are presented in Table 1. The delay lengths are prime numbers distributed logarithmically. In all configurations, the total number of modes is $90008700<\mathcal{M}<$ 9000, which ensures us to rule out artifacts related to low modal density [16].
Table 1: Delay-line lengths (in samples) for each size $N$ of the analyzed FDNs. The lengths are logarithmically distributed prime numbers.
@@
During optimization, we used a sampling rate of $f_{\textrm{s}}=48$ kHz and $M=480000$ frequency points evenly distributed over $[0,\pi]$. The batch size was $\mu=2000$, requiring 240 steps to complete an epoch. For training, we used $80\%$ of the frequency points, while the remaining $20\%$ were used for validation. We employed Adam optimizer [31] with learning rate of $\eta=10^{-3}$. Regarding the gain-per-sample, we found that $\gamma=0.9999$ leads to stable training and fast convergence. This choice implies $T_{\textrm{60}}=1.439\,$s and satisfies (15).
The configurations under analysis include the differentiable FDN optimized according to our prior work (DAFx23) [9], which employs an orthogonal feedback matrix, as well as the one optimized with the proposed method (DiffFDN-O) that also features an orthogonal feedback matrix, and two variations of the latter: one with an optimizable Householder matrix (DiffFDN-HH) and another with an optimizable scattering matrix (DiffFDN-SCAT).
@@ 5.2 Computational complexity
The presented optimization utilizes a sparse set of frequency sampling points at each step. This choice is driven by the aim to decrease the overall number of operations, providing a notable reduction when compared to the baseline method [9]. During the forward pass, the computational complexity needed to calculate the FDN transfer function at a single frequency point in (11) is $\mathcal{O}(N^{3})$. For each epoch, the proposed method (DiffFDN-O) computes the transfer function at $M$ points, where $M$ corresponds to the dataset size. In contrast, the baseline approach (DAFx23) requires calculating a similar number of frequency points at each training step, leading to a linear increase in the number of operations as the dataset size grows, which in [9] it is represented by different values of $M$ (e.g., 256). This is because the number of epochs required for convergence in the baseline is close to that of the presented method. Additionally, in the forward pass of the baseline, the DFT is necessary due to the temporal loss, introducing additional operations. By employing the Householder matrix, we can significantly decrease the computational complexity of (11) to $\mathcal{O}(N)$. Introducing the DFT would raise this complexity to $\mathcal{O}(N\text{log}_{2}(N))$.
@@
DiffFDN-O 208 324 448
DiffFDN-HH 200 300 400
DiffFDN-SCAT 288 480 704
@@ 5.3 Modal excitation distribution
Table 3: Standard deviation of the modal excitation of DiffFDN-O before (Init) and after optimization (Optim) at different FDN sizes.
@@
To assess the impact of the changes to our previous optimization framework [9]—particularly the addition of the sparsity loss function and sparse frequency sampling—we ran the same tests with the original optimization framework for comparison. Each set of parameters used at initialization is shared between the two methods. We observed comparable convergence rates for both optimization methods. Consequently, we allowed each training run to complete 20 epochs before halting it and extracting the output parameters for evaluation.
@@
Table 4 also shows the results of a similar test run on DiffFDN-SCAT and DiffFDN-HH, both of size $N=4$. Due to a different parameterization of the feedback matrix, the distributions of the initial parameters are different from both DAFx23 and DiffFDN-O. After optimization, the standard deviation reduces in both cases. Among all tested configurations, the DiffFDN-HH case exhibited the smallest improvement, likely due to its more restrictive parameterization. The DiffFDN-SCAT case also showed a lower degree of improvement compared to DiffFDN-O. This could be attributed to the higher complexity of the problem, suggesting potential enhancement through fine-tuning the optimization framework.
The histograms in Fig. 6 illustrate the distribution of modal excitation at the start (Init) and end of the optimization process (Optim) for a single instance of DiffFDN-O, across all analyzed sizes. The modal excitation values have been centered around 0 dB. The optimization reduces the loudest modes, leading to a shift toward a narrower overall excitation distribution. Figure 7 shows the effect on their magnitude spectrum. Similarly to Fig. 2, only a narrow section of the spectrum is displayed. The optimized responses show a more uniform behavior. While resonances are inherent to the FDN structure, their prominence has been reduced and more evenly distributed. The effect of these changes on perceived coloration is further evaluated through a subjective test.
Figure 6: Distribution of the modal excitation of an FDN with sizes $N=4,6,\,\text{and}\,\,8$ at the beginning (Init) and at the end of optimization (Optim), which has led to a decrease of the loudest modal excitation and to a narrowing of the distribution.
Figure 7: Section of the magnitude response of the FDNs in Figure 6. The responses at initialization (Init) and at the end of optimization (Optim) have been shifted apart by 50 dB for better clarity. During optimization, the resonances have been reduced and more evenly distributed.
Table 4: Standard deviation of the modal excitation before (Init) and after optimization (Optim) of different frameworks, each employing an FDN of size N=4 and identical set of delays.
@@
DiffFDN-O 7.8346 4.4518
DiffFDN-HH 8.5641 7.8765
DiffFDN-SCAT 7.6200 5.8822
@@ 6.1 Listening test procedure
Figure 8: Listening test results on the analyzed reverberation conditions: (a) lossless, (b) homogeneous decay, and (c) frequency-dependent decay. The three DiffFDN conditions correspond to FDN configurations trained using the proposed optimization framework.
@@
In the part testing the lossless condition, we evaluated late non-decaying reverb. The reference for this part was a 3-second segment of white Gaussian noise, known to be highly colorless. The IRs of the FDNs were extracted from the mixing time, thereby eliminating the initial echo buildup from the evaluation. The section testing homogeneous decay, focused on frequency-independently decaying IRs, with a $T_{60}$ value set at 2 s for all frequencies. A corresponding exponential decaying envelope with the same $T_{60}$ value was applied to the reference white Gaussian noise. The part testing frequency-dependent decay involved decaying IRs with a fixed frequency dependent $T_{60}(\omega)$ across conditions. The reference in this case was the IRs of an FDN of size $N=64$, ensuring a sufficiently smooth response. The frequency-dependent attenuation filter was designed following the steps outlined in Section 4.1.
The test evaluated the coloration of the standard differentiable FDN with orthogonal feedback matrix (DiffFDN-O) proposed in this paper, the DiffFDN with scattering matrix (DiffFDN-SCAT), and that with optimizable Householder matrix (DiffFDN-HH). Each configuration was tested on a separate page where the number and lengths of the delays were constant, and only the feedback matrix and the input and output gains were altered. The FDN implementation of the Schroeder series allpass reverberator (SH) was the anchor, whereas the random orthogonal feedback matrix (RO) acted as a baseline. For the RO condition, the initial values of optimization of DiffFDN-O were used. The direct gain $d$ was set to zero in all cases. We excluded the DiffFDN trained with the framework presented in [9] since it would constitute a condition perceptually nearly identical to DiffFDN-O, and as such might skew the results. The primary contribution of this work is focused on improving training efficiency. In total, there were 18 listening test pages with six stimuli each. The loudness of each IR was normalized to ensure consistency across conditions.
@@
The $p$-values for most paired conditions indicate significant differences among all pairs of results. However, there are a few exceptions observed in the distribution of DiffFDN-O and DiffFDN-HH. In these cases, the $p$-values exceed the alpha level across all conditions with decaying modes, suggesting a convergence in the scores assigned to the two conditions. Moreover, DiffFDN-O and DiffFDN-HH reject the null hypothesis when paired with DiffFDN-SCAT in the lossy cases for $N=8$, suggesting that for larger FDN sizes, the performance improvement provided by DiffFDN-SCAT diminishes.
The IRs of the optimized FDNs always show a median greater than the initial condition (RO), proving the effectiveness of the optimization. In the late reverb analysis, Fig. 8 (a), the DiffFDN-O has median ratings of 74, 82.5, and 70, respectively, for increasing FDN size. Only for $N=$ 8, the FDN with scattering matrix received ratings higher than DiffFDN-O with a median of 76.
In the homogeneous decay and frequency-dependent decay tests the full IRs are taken into account. The scores given to the DiffFDN-SCAT condition suggest that the scattering matrix contributes to the smoothness of the early reflections as well as the tail. In general, spectral flatness is associated with temporal smoothness. The scattering feedback matrix increases echo density during early reflections, potentially suggesting a less colored response, as shown in Fig. 8 (b) and (c). This effect becomes more noticeable for smaller FDNs, and for $N=8$, the medians of the results begin to converge among optimized FDNs.
The confidence intervals for the results of all tests are relatively large. Although the objective results in Table 4 indicate a limited reduction in modal excitation standard deviation, the optimized Householder matrix was rated relatively close to the DiffFDN-O in the listening test. The distribution of modal excitation is one quantity that can aid in deducing the coloration of the reverberator. However, this is a much more complicated concept, as evidenced by the results presented in the listening test.